OpenStudy (anonymous):

log x radical x^2+1 i have to write as a sum

OpenStudy (anonymous):

\[here's how \it looks, logx \sqrt{x ^{2}+1}\]

OpenStudy (anonymous):

Hints: \(\Large \log(\color{red}a \color{blue}b) = \log (\color{red} a ) +\log( \color{blue} b)\) \(\Large \log( a^{\color{green} n})= \color{green} n \log (a)\)